# SS-GEN: A Social Story Generation Framework with Large Language Models

### v2

@@ Abstract
Children with Autism Spectrum Disorder (ASD) often misunderstand social situations and struggle to participate in daily routines. Social Stories™ [^1] As the social-skill intervention tool, Social Stories™ are widely used to specifically help individuals with ASD understand and navigate specific social situations. When capitalized, a Social Story refers to a Social Story™ following the criteria in 10.2 Tutorial (Gray 2010). are traditionally crafted by psychology experts under strict constraints to address these challenges but are costly and limited in diversity. As Large Language Models (LLMs) advance, there’s an opportunity to develop more automated, affordable, and accessible methods to generate Social Stories in real-time with broad coverage. However, adapting LLMs to meet the unique and strict constraints of Social Stories is a challenging issue. To this end, we propose SS-GEN, a Social Story GENeration framework with LLMs. Firstly, we develop a constraint-driven sophisticated strategy named StarSow to hierarchically prompt LLMs to generate Social Stories at scale, followed by rigorous human filtering to build a high-quality dataset. Additionally, we introduce quality assessment criteria to evaluate the effectiveness of these generated stories. Considering that powerful closed-source large models require very complex instructions and expensive API fees, we finally fine-tune smaller language models with our curated high-quality dataset, achieving comparable results at lower costs and with simpler instruction and deployment. This work marks a significant step in leveraging AI to personalize Social Stories cost-effectively for autistic children at scale, which we hope can encourage future research. The prompt, code and data will release in the Technical Appendix and Code & Data Appendix at https://github.com/MIMIFY/SS-GEN.
@@ Introduction
Social Story, pioneered by Gray and Garand (1993), helps children with autism [^2] Autism Spectrum Disorder (ASD) is a lifelong neurodevelopmental disorder characterized by impairments in social interaction, communication, and restricted or repetitive behaviors, affecting millions of children (approximately 1/100) worldwide with the number continuously growing (Zeidan et al. 2022). define the social context of anticipated or experienced situations, putting them into perspective and then developing statements on how they feel and what actions to take in response, thereby positively improving their social understanding and behavioral outcomes. For example, the story in Figure 1 helps autistic children understand and manage the stress and anxiety triggered by mistakes.
@@
In this work, we introduce SS-GEN, a Social Story GENeration framework based on their unique constraints. (1) Initially, we develop a constraint-driven strategy named StarSow (Figure 2) to construct a dataset of 5K+ Social Stories by prompting advanced LLMs (e.g., GPT-4o) through a breadth-first, hierarchical synthesis from a seed set of manually-written stories, followed by rigorous human filtering. The StarSow is similar to writing a book or sowing a star-tree. Concretely, we first generate chapters to cover various story themes (roots), then create story titles (branches) for each theme as much as possible, and last generate story content (star fruits) cautiously that meets specific constraints given the title and the theme. The analysis shows that our dataset is more diverse, flexible, and effective, making it acceptable to human evaluators. (2) Additionally, to verify the effectiveness of generated Social Stories and ensure their adherence to the strict constraints above, we propose the quality assessment criteria which is employed in both human and GPT evaluations. (3) Finally, we train smaller language models on our crafted dataset, considering that powerful closed-source LLMs require very complex instructions and expensive API fees. Empirical results show that our dataset significantly improves LLMs performance on SS-GEN, at lower costs and with simpler deployment, making automated Social Stories more accessible and affordable to the autism community.
In summary, to the best of our knowledge, we are the first to propose SS-GEN task and framework, which aims to make Social Stories more accessible, affordable, and of high quality for their use as a social-skill intervention tool within the autism community. Our framework includes:
@@
Q4: Does the Social Story use exact vocabulary? The Social Story should choose the most comfortable and accurate vocabulary for the audience. Firstly, use positive verbs while also being mindful of the varying implications of verbs. Avoid using terms that are likely to evoke strong emotional feelings such as ”shouldn’t”, ”must”, ”supposed to” and so on.
@@
Gardening Work. To encourage diversity, a new chapter is added to the Social Story dataset only when its ROUGE-L similarity with any existing chapter is less than 0.7. The same applies to a new Social Story title. When generating the Social Story content based on the title goal, we filter out invalid outputs that do not meet the strict constraints of the Social Stories (e.g., using the narrative ’you’, or negative words that usually cause anxious or resistant feelings). After manual screening and refinement, only 74.3% of the data was deemed valid to construct the final dataset.
@@
Figure 3: The top 20 most common root verbs (inner circle) and their top 5 direct noun objects (outer circle) in the generated Social Story titles.
@@
To further analyze what types of intervention goals are generated and how diverse they are, we identify the verb-noun structure in the generated Social Story titles. Through the Berkeley Neural Parser [^4] https://parser.kitaev.io/ (Kitaev and Klein 2018; Kitaev, Cao, and Klein 2019), we parse the titles and then extract the verb closest to the root along with its first direct noun object. Figure 3 shows the top 20 most common root verbs and their top 5 direct noun objects, which can represent common topics within the dataset. The figure shows these intervention goals exhibit a wide range of intents and textual formats.
@@
In this section, we conduct extensive experiments by training and testing smaller, cost-effective language models using our crafted dataset for SS-GEN, considering that powerful closed-source LLMs require very complex instructions and expensive API fees.
@@
The code and data, StarSow prompts and experimental hyperparameters have been uploaded in the supplementary Technical Appendix, Code & Data Appendix and will be released.
Title-to-Story for SS-GEN Develop a concise, clear, straightforward, positive and supportive Social Story titled ”{title}” for children and teens with autism, 200-300 words, that promotes their social understanding and boosts their participation in daily activities, fostering independence and confidence.
@@
Figure 6: GPT-4 evaluation results of various models on five dimensions on Zero-Shot and SFT scenarios for SS-GEN: Coherence ($\mathcal{CH.}$), Descriptiveness ($\mathcal{DC.}$), Empathy ($\mathcal{EM.}$), Grammaticality ($\mathcal{GA.}$), and Relevance ($\mathcal{RE.}$).
@@
For GPT-4 evaluation, we apply the Coherence ($\mathcal{CH.}$), Descriptiveness ($\mathcal{DC.}$), Empathy ($\mathcal{EM.}$) scores to separately evaluate whether the Social Story meets the checklist requirements for structural clarity, descriptive orientation, situational safety as outlined in the criteria. Additionally, for the text itself, we use the Grammaticality ($\mathcal{GA.}$) score to judge grammatical correctness, and the Relevance ($\mathcal{RE.}$) score to assess the consistency between the generated story content and the given title prompt. All scores range from 1 to 5, with 5 representing the best.
For human evaluation, we invite three psychology students who received comprehensive training on the quality assessment criteria guidelines before their participation in the evaluation. Additionally, we develop a user-friendly website shown in the supplementary to standardize expert evaluations, improve efficiency, and enhance the horizontal comparison of text generated by different models. Based on the criteria, experts use this website to score Q1, Q2, Q3, Q4 of the structural clarity ($\mathbb{SC.}$) on a scale of 1 to 5, and to provide ’Yes/No’ for Q1 of the descriptive orientation ($\mathbb{DO.}$) and Q1(A), Q1(B), Q2, Q3, Q4 of the situational safety ($\mathbb{SS.}$). Finally, we report the average score of all $\mathbb{SC.}$ sub-items, as well as the proportion of qualified $\mathbb{SS.}$ and $\mathbb{DO.}$, where the “qualified” means all sub-items are rated as “Yes”.
@@
We evaluate mainstream models of various scales for generating Social Stories in Zero-Shot and SFT scenarios. Table 4 presents the average results of three inferences on the test set using traditional metrics. Figure 6 shows GPT-4 evaluations of 200 randomly selected stories from the test set, focusing on Coherence ($\mathcal{CH.}$), Descriptiveness ($\mathcal{DC.}$), Empathy ($\mathcal{EM.}$), Grammaticality ($\mathcal{GA.}$) and Relevance ($\mathcal{RE.}$). In Table 5, we select the top 2B and 7B models on both traditional and GPT-4 evaluations for further expert assessment, along with 50 randomly selected stories from the test set for the expert assessment. The case study have been uploaded in the supplementary Technical Appendix.

### v3

@@ Abstract
Children with Autism Spectrum Disorder (ASD) often misunderstand social situations and struggle to participate in daily routines. Social Stories™ are traditionally crafted by psychology experts under strict constraints to address these challenges but are costly and limited in diversity. As Large Language Models (LLMs) advance, there’s an opportunity to develop more automated, affordable, and accessible methods to generate Social Stories in real-time with broad coverage. However, adapting LLMs to meet the unique and strict constraints of Social Stories is a challenging issue. To this end, we propose SS-GEN, a Social Story GENeration framework with LLMs. Firstly, we develop a constraint-driven sophisticated strategy named StarSow to hierarchically prompt LLMs to generate Social Stories at scale, followed by rigorous human filtering to build a high-quality dataset. Additionally, we introduce quality assessment criteria to evaluate the effectiveness of these generated stories. Considering that powerful closed-source large models require very complex instructions and expensive API fees, we finally fine-tune smaller language models with our curated high-quality dataset, achieving comparable results at lower costs and with simpler instruction and deployment. This work marks a significant step in leveraging AI to personalize Social Stories cost-effectively for autistic children at scale, which we hope can encourage future research on special groups.
@@ Introduction
Social Story™ [^1] As the social-skill intervention tool, Social Stories™ are widely used to specifically help individuals with ASD understand and navigate specific social situations. When capitalized, a Social Story refers to a Social Story™ following the criteria in 10.2 Tutorial (Gray 2010)., pioneered by Gray and Garand (1993), helps children with autism [^2] Autism Spectrum Disorder (ASD) is a lifelong neurodevelopmental disorder characterized by impairments in social interaction, communication, and restricted or repetitive behaviors, affecting millions of children (approximately 1/100) worldwide with the number continuously growing (Zeidan et al. 2022). define the social context of anticipated or experienced situations, putting them into perspective and then developing statements on how they feel and what actions to take in response, thereby positively improving their social understanding and behavioral outcomes. For example, the story in Figure 1 helps autistic children understand and manage the stress and anxiety triggered by mistakes.
@@
In this work, we introduce SS-GEN, a Social Story GENeration framework based on their unique constraints. (1) Initially, we develop a constraint-driven strategy named StarSow (Figure 2) to construct a dataset of 5K+ Social Stories by prompting advanced LLMs (e.g., GPT-4o) through a breadth-first, hierarchical synthesis from a seed set of manually-written stories, followed by rigorous human filtering. The StarSow is similar to writing a book or sowing a star-tree. Concretely, we first generate chapters to cover various story themes (roots), then create story titles (branches) for each theme as much as possible, and last generate story content (star fruits) cautiously that meets specific constraints given the title and the theme. The analysis shows that our dataset is more diverse, flexible, and effective, making it acceptable to human evaluators. (2) Additionally, to verify the effectiveness of generated Social Stories and ensure their adherence to the strict constraints above, we propose the quality assessment criteria which is employed in both human and GPT evaluations. (3) Finally, we train smaller language models on our crafted dataset, considering that powerful closed-source LLMs require very complex instructions and expensive API fees. Empirical results show that our dataset significantly improves LLMs performance on SS-GEN, at lower costs and with simpler instructions.
In summary, to the best of our knowledge, we are the first to propose SS-GEN task and framework, which aims to make Social Stories automated, affordable, more accessible and of high quality for their use as a social-skill intervention tool within the autism community. Our framework includes:
@@
Q4: Does the Social Story use exact vocabulary? The Social Story should choose the most comfortable and accurate vocabulary for the audience. Firstly, use positive verbs while also being mindful of the varying implications of verbs. Avoid using terms that are likely to evoke strong emotional feelings such as “shouldn’t”, “must”, “supposed to” and so on.
@@
Gardening Work. To encourage diversity, a new chapter is added to the Social Story dataset only when its ROUGE-L similarity with any existing chapter is less than 0.7. The same applies to a new Social Story title. When generating the Social Story content based on the title goal, we filter out invalid outputs that do not meet the strict constraints of the Social Stories (e.g., using the narrative “you”, or negative words that usually cause anxious or resistant feelings). After manual screening and refinement, only 74.3% of the data was deemed valid to construct the final dataset.
@@
Figure 3: The top 20 most common root verbs (inner circle) and their top 3 direct noun objects (outer circle) in the generated Social Story titles.
@@
To further analyze what types of intervention goals are generated and how diverse they are, we identify the verb-noun structure in the generated Social Story titles. Through the Berkeley Neural Parser [^4] https://parser.kitaev.io/ (Kitaev and Klein 2018; Kitaev, Cao, and Klein 2019), we parse the titles and then extract the verb closest to the root along with its first direct noun object. Figure 3 shows the top 20 most common root verbs and their top 3 direct noun objects, which can represent common topics within the dataset. The figure shows these intervention goals exhibit a wide range of intents and textual formats.
@@
In this section, we conduct extensive experiments [^5] The code, prompt, data and technical appendix are available at https://github.com/MIMIFY/SS-GEN to evaluate the performance of smaller, cost-effective language models which are trained and tested using our curated dataset for SS-GEN, considering that powerful closed-source LLMs require very complex instructions and expensive API fees.
@@
Title-to-Story for SS-GEN Develop a concise, clear, straightforward, positive and supportive Social Story titled “{title}” for children and teens with autism, 200-300 words, that promotes their social understanding and boosts their participation in daily activities, fostering independence and confidence.
@@
Figure 6: GPT-4 evaluation results of various models on five dimensions on Zero-Shot (blue) and SFT (orange) scenarios for SS-GEN: Coherence ($\mathcal{CH.}$), Descriptiveness ($\mathcal{DC.}$), Empathy ($\mathcal{EM.}$), Grammaticality ($\mathcal{GA.}$), and Relevance ($\mathcal{RE.}$).
@@
For GPT-4 evaluation, we apply the Coherence ($\mathcal{CH.}$), Descriptiveness ($\mathcal{DC.}$), Empathy ($\mathcal{EM.}$) scores to separately evaluate whether the Social Story meets the checklist requirements for structural clarity, descriptive orientation, situational safety as outlined in the established criteria for quality assessment. Additionally, for the text itself, we use the Grammaticality ($\mathcal{GA.}$) score to determine grammatical correctness, fluency, and adherence to language standards. We also use the Relevance ($\mathcal{RE.}$) score to assess the consistency and alignment between the generated story content and the given title prompt. All scores range from 1 to 5, with 5 representing the highest level of quality, clarity, and adherence to the evaluation standards.
For human evaluation, we invite three psychology students who received comprehensive training on the quality assessment criteria guidelines before their participation in the evaluation. Additionally, we develop a user-friendly website shown in the supplementary to standardize expert evaluations, improve efficiency, and enhance the horizontal comparison of text generated by different models. Based on the criteria, experts use this website to score Q1, Q2, Q3, Q4 of the structural clarity ($\mathbb{SC.}$) on a scale of 1 to 5, and to provide “Yes/No” for Q1 of the descriptive orientation ($\mathbb{DO.}$) and Q1(A), Q1(B), Q2, Q3, Q4 of the situational safety ($\mathbb{SS.}$). Finally, we report the average score of all $\mathbb{SC.}$ sub-items, as well as the proportion of qualified $\mathbb{SS.}$ and $\mathbb{DO.}$, where the “qualified” means all sub-items are rated as “Yes”.
@@
We evaluate mainstream models of various scales for generating Social Stories in Zero-Shot and SFT scenarios. Table 4 presents the average results of three inferences on the test set using traditional metrics. Figure 6 shows GPT-4 evaluations of 200 randomly selected stories from the test set, focusing on Coherence ($\mathcal{CH.}$), Descriptiveness ($\mathcal{DC.}$), Empathy ($\mathcal{EM.}$), Grammaticality ($\mathcal{GA.}$) and Relevance ($\mathcal{RE.}$). In Table 5, we select the top 2B and 7B models on both traditional and GPT-4 evaluations for further expert assessment, along with 50 randomly selected stories from the test set for the expert assessment. We present the analysis of the case study in the technical appendix.
@@ Conclusion
Ethical Statement
The generation of Social Stories entails a range of ethical considerations and limitations that require careful attention to ensure their responsible application, particularly in sensitive contexts involving neurodiverse children. The primary ethical issues are outlined as follows. (1) Bias and Stereotypes: training data may carry social biases although we have rigorous human review to ensure quality. (2) Caregiver supervision: AI-drafted stories need caregiver’s refinement to suit each child’s unique experiences, ensuring ethical and emotional suitability. (3) Ethical risks: limited transparency in generation may affect long-term predictability. Expert feedback and heuristic controls are advised to mitigate emotional insensitivity risks.
Acknowledgments
This work was partly supported by the National Key Research and Development Program of China under Grant 2024YFE0202900; the National Natural Science Foundation of China under Grant (62436001, 62176020); the Joint Foundation of the Ministry of Education for Innovation team (8091B042235); the Fundamental Research Funds for the Central Universities (2019JBZ110) ; and the State Key Laboratory of Rail Traffic Control and Safety (Contract No. RCS2023K006), Beijing Jiaotong University. We sincerely thank our collaborators for their invaluable contributions, including insightful feedback and support in refining the paper and addressing key challenges. Their dedication and expertise greatly enhanced the quality and impact of this work. We thank the anonymous reviewers for carefully reading our paper and their insightful comments and suggestions.